Hrec has two actuator models for each mirror NE,WE,BS: one model for LN1, one model for LN2.
When the flag V1:SAT_NE_LN2_P2 is 1, Hrec uses the LN2 model for NE
When the flag V1:SAT_WE_LN2_P2 is 1, Hrec uses the LN2 model for WE
When the flag V1:SAT_BS_LN2_P2 is 1, Hrec uses the LN2 model for BS
If those flags are missing, Hrec considers by default that ITF is in LN2.
If those flags are at 0, Hrec considers that we are in LN1.
Since the relock of this afternoon, the ITF was in LN2 but the flags were kept at 0, thus Hrec was using LN1 models for NE,WE,BS actuators.
That's why the sensitivity was more noisy (BNS range around 48 Mpc instead of 55 Mpc) and the adjustment of optical response was wrong (DCP frequency around 260 Hz instead of 180 Hz).
Waiting for the SAT flags to be put back to the correct values of 1, I have restarted Hrec arounf 21h45 UTC after having commented the lines SWITCH_LN2 and SET_ACTUATOR in the configuration Hrec_actuators.cfg (by default Hrec considers that ITF is in LN2). So, we are back to 55 Mpc.
Next time suspension's people wants to do an April fool with the coil drivers flags, they should tell us before the end of the day. We loose 6 hours in finding the origin of the wrong h(t) reconstruction.